# IDOL: Instant Photorealistic 3D Human Creation from a Single Image

### v1

@@ 2 Related Work
High-precision 3D human models typically rely on scanning or multi-view camera acquisition systems. Scan-based datasets, such as THuman2.0 [64], Twindom [21], and 2K2K [14], provide high-fidelity 3D geometries but are limited by a small number of subjects, simple standing poses, and non-human artifacts (e.g., 2K2K). In contrast, multi-view acquisition systems [46, 37, 66, 34, 4, 8, 27, 9, 59] facilitate the collection of larger datasets with more flexible actions. However, these datasets often face challenges such as biased indoor lighting, fixed viewpoints, and limited scale. Large-scale synthetic and real datasets for generic objects, including Objaverse and MVImgNet [11, 12, 65], address open-world reconstruction but lack specificity for human models. To overcome these limitations, the proposed HuGe100K significantly scales up dataset size, increasing subject diversity by over 100 times compared to previous datasets. It emphasizes diversity in pose, shape, viewpoint, and clothing, paving the way for large-scale model training. We compare existing datasets with ours in Tab. 1.
@@ 4.1 Animatable Human Representation
To estimate the skinning weights for each $\mathcal{G}_{k}$, we compute the skinning weight field for body parts using a pre-defined low-resolution volume[7]. For the smaller regions, such as the hands and face, which exhibit minimal topology changes, their skinning weights are calculated by interpolating from the SMPL-X template using barycentric coordinates. This approach supports modeling large-scale topology changes among different identities, including substantial changes introduced by clothing. In addition, it helps stabilize the convergence for areas like the fingers and face, which experience fewer topological changes compared to the SMPL-X template.
@@ 4.2 Network Structure
As illustrated in Fig. 4, IDOL is a full-differentiable transformer-based framework for reconstructing animatable 3D human from a single image.
@@ 5.2 Comparison of Reconstruction Model
For the encoder, we employed the pre-trained Sapiens-1B model[29] and kept its weights frozen during training. We then defined a transformer-based framework with a parameter size of $0.5$ billion to perform feature fusion. Additionally, we densify the SMPL-X vertex set by sampling approximately 200,000 vertices to map attributes and represent the 3D human model effectively.
@@ 6 Conclusions and Limitations
However, there are limitations. Due to the constraints of the video model used, we can only synthesize single-frame images from fixed viewpoints. Future work could consider generating longer motion sequences. Additionally, our dataset lacks the generation of data involving multiple people interacting or people interacting with objects. Our method also struggles with handling half-body inputs. Future efforts could focus on enhancing data generation strategies to improve model performance.

### v2

@@ 2 Related Work
Recent advancements in 3D avatar reconstruction [14, 90, 76, 22, 15, 13, 33] and representation learning [68, 16, 92, 91] have greatly improved human modeling. Most relevant to our work are recent efforts on 3D human datasets and single-image human reconstruction.
@@
High-precision 3D human models typically rely on scanning or multi-view camera acquisition systems. Scan-based datasets, such as THuman2.0 [75], Twindom [28], and 2K2K [19], provide high-fidelity 3D geometries but are limited by a small number of subjects, simple standing poses, and non-human artifacts (e.g., 2K2K). In contrast, multi-view acquisition systems [55, 45, 78, 42, 4, 8, 35, 9, 70, 23, 93] facilitate the collection of larger datasets with more flexible actions. However, these datasets often face challenges such as biased indoor lighting, fixed viewpoints, and limited scale. Large-scale synthetic and real datasets for generic objects, including Objaverse and MVImgNet [11, 12, 77], address open-world reconstruction but lack specificity for human models. To overcome these limitations, the proposed HuGe100K significantly scales up dataset size, increasing subject diversity by over 100 times compared to previous datasets. It emphasizes diversity in pose, shape, viewpoint, and clothing, paving the way for large-scale model training. We compare existing datasets with ours in Tab. 1.
@@ 4.1 Animatable Human Representation
To estimate the skinning weights for each $\mathcal{G}_{k}$, we first compute a body-part skinning field using a pre-defined low-resolution volume [7]. For smaller regions like the hands and face, which undergo minimal topological variation, weights are interpolated from the SMPL-X template via barycentric coordinates. This strategy enables modeling large topology changes across identities (e.g., from clothing) while stabilizing convergence in less variable areas such as fingers and facial regions.
@@ 4.2 Network Structure
As illustrated in Fig. 4, IDOL is a full-differentiable framework for reconstructing animatable 3D human.
@@ 5.2 Comparison of Reconstruction Model
For the encoder, we employ the pre-trained Sapiens-1B model [37] and keep its weights frozen during training. We then define a transformer-based framework with a parameter size of $0.5B$ to perform feature fusion. Additionally, we densify the SMPL-X vertex set by sampling approximately $200,000$ vertices to map attributes and represent the 3D human model effectively.
@@ 6 Conclusions and Limitations
However, there are limitations. Due to the constraints of the video model used, we can only synthesize single-frame images from fixed viewpoints. Future work could consider generating longer motion sequences. The focus on body reconstruction and animation leaves facial optimization secondary, and the architecture lacks specific design for facial identity or expression. Additionally, handling half-body inputs remains challenging, and improvements in data generation strategies are needed to enhance performance.
Acknowledgments
This work was partly supported by the National Key Research and Development Program of China (2022YFF0902400), the Shenzhen Science and Technology Program (JCYJ20220818101014030), and the research fund of Tsinghua University - Tencent Joint Laboratory for Internet Innovation Technology.
@@ Appendix C Experiment
More Qualitative Comparisons.
We show additional visual comparisons in Fig. 14 and Fig. 15. We compare with the reported results by Weng et al. [66] and AlBahar et al. [2].
Effect of the SMPL-X Parameters on Reconstruction.
Although the reconstruction quality remains good with imperfect SMPL-X input, errors such as leaning or bent shapes can occur due to inaccurate pose parameters, as shown in Fig. 6 of the main content. This occurs because the avatar is re-posed based on the estimated SMPL-X parameters. Fig.16b demonstrates that providing accurate pose information resolves this issue.
Dataset WE (×$\times 10^{-3}$) $\downarrow$
THuman2.1 5.38
HuGe100K (MVChamp) 7.33
Zero123 10.51
Table 3: Warping Error (WE) comparison across datasets and multi-view synthesis methods, evaluating 3D consistency.
Method MSE $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$
SIFU 0.032 15.054 1.303
GTA 0.035 14.833 1.340
Ours 0.023 16.688 1.171
Table 4: Quantitative comparison on the 2K2K dataset.
Figure 16: (a) Comparison with Zero123. (b) Original results with leaning/bent poses due to inaccurate SMPL-X, and corrected results with refined SMPL-X. (c) The results on 2K2K. (d) Challenges in large pose and loose cloth animation. (e) 3D animating framework using TRELLIS for image-to-3D and Make-It-Animatable for rigging and animation.
Experimental Comparison with Other Multi-View Image Generation Models.
Here, we compare MVChamp with traditional multi-view image generation models based on text-to-image synthesis, specifically Zero123, in the context of human multi-view generation. We compare WE [46] in Tab. 3, evaluating 50 random cases. The THuman2.1 dataset serves as the upper bound, and HuGe100K shows comparable results; Regarding multi-view generation, MVChamp outperforms Zero123 by 30.1% in 3D consistency. Zero123 generates one novel view at a time, causing multi-view inconsistency. In contrast, MVChamp generates 24 views per batch, ensuring consistency. It also benefits from redundant human priors from dance videos and provides more accurate pose control, enabling well-aligned SMPL-X parameters. See Fig. 16a for a visual comparison.
Additional Cases for Evaluating Generalization to Complex Poses and Loose Clothing.
Fig. 16d demonstrates IDOL’s capability to handle complex poses and loose clothing. This is made possible by our novel architecture, which extracts global features using Sapiens and the diverse HuGe100K dataset. While loose clothing presents challenges due to significant deviation from the body, HuGe100K provides numerous examples, allowing IDOL to recover animatable 3D avatars effectively and reduce issues like tearing in animations, especially in areas such as skirts. For more examples and animations, please refer to the introduction video (38s-54s).
Comparison with 3D Animating Methods.
Classical animation methods typically involve image-to-3D conversion, rigging, and animation. Fig.16e demonstrates this pipeline using TRELLIS[69] and Make-It-Animatable [17], which struggles with topology changes, such as detaching the hand from the waist, resulting in artifacts. In contrast, our approach (left) handles these transitions naturally.
Evaluation on Additional 3D Datasets.
We performed the suggested evaluation on 2K2K using the same settings as in the paper. The quantitative and qualitative results are presented in Tab.4 and Fig.16c, offering valuable insights into IDOL’s generalizability.